# Johan Håstad

### v2

@@ 4 The switching lemma
Both $\tau$ and $\tau^{*}$ are essentially almost complete matchings. We do not allow $\tau$ to be lopsided but this only reduces the number of possibilities, by Lemma 3.7, by a factor $1+o(1)$. It is the case that $\tau^{*}$ is only slightly more general but we only need an upper bound on the number of possibilities. On the other hand if $U^{*}$ does not contain the designated survivor there are two mini-squares in $\tau^{*}$ where the sum is not zero. This is the designated survivor and its replacement in $U^{*}$ in the same super-square. The number of $\tau^{*}$ with these fixed sums is, however, smaller by Lemma 3.8. There is no difference between $B$ and $B^{*}$ and thus the number of alternatives is the same. We conclude that if we denote the number of different $\tau,U$ and $B$ by $N$, then the number of triples $\tau^{*},U^{*}$, and $B^{*}$ is bounded by $\Delta N(1+o(1))$.
@@
The big difference is in the probability of picking the restriction $\pi_{2}$ and $\pi_{2}^{*}$. Let $D$ denote the total number of possible pairs of mini-squares. We have that $D=\Delta^{2}(n/T)^{2}(2+o(1))$. This follow as there are $\Delta(n/T)^{2}$ mini-squares. Most of them (except at the perimeter) have $4\Delta$ possible partners and hence the number of pairs is $(1+o(1))\Delta(n/T)^{2}4\Delta/2$. We have the following lemma.
Lemma 4.9
@@
Assuming $\Delta=\omega(\log n)$ then $Pr[\pi_{2}]\leq 2^{o(s)}(k/D)^{s}Pr[\pi_{2}^{*}]$ which in its turn is bounded by $2^{O(s)}(\log n/\Delta^{2})^{s}Pr[\pi_{2}^{*}]$.
@@
Proof:  The big difference between the two is that $\pi_{2}$ is of size $k$ while $\pi_{2}^{*}$ is of size only $k-s$. The former is also not unbalanced, but as this, by Lemma 3.9, only changes the probability by a factor $(1+o(1))$ and we can let this factor be part of the error term. The way we define the probability space of partial matchings is slightly unusual and we need to be careful. There are $k!$ different orders to pick $\pi_{2}$. The probability of picking $\pi_{2}$ in a particular order $\gamma$ is $\prod_{i=1}^{k}t_{i,\gamma}^{-1}$ where $t_{i,\gamma}$ is the number of pairs of mini-squares available at stage $i$ assuming the order $\gamma$.
Similarly the probability of $\pi_{2}^{*}$ can be written as
@@
where $\gamma^{*}$ determines the order in which its $k-s$ pairs are added. Each picked pair of mini-squares eliminates at most $8\Delta$ possible other pairs and hence each $t_{i,\gamma}$ and $t_{i^{*},\gamma^{*}}$ is at least $D-8k\Delta$. The product giving the probabilities do have many factors so we need more precise information.
Say that $\gamma$ and $\gamma^{*}$ are associated (written as $\gamma\sim\gamma^{*}$) if the $k-s$ common elements are in the same order. In such a situation for each $i^{*}$ there is an $i$ corresponding to the same element such that $t_{i^{*},\gamma^{*}}\geq t_{i,\gamma}\geq t_{i^{*},\gamma^{*}}-8s\Delta$. Restricting the product to these paired indices and using $t_{i,\gamma}\geq D/2$ we have
@@
$(1-16s\Delta/D)^{k}=2^{-\Theta((s\log n)/\Delta)}=2^{o(s)},$
@@
giving the bound of the lemma. The second bound is obtained by using the definitions of $k$ and $D$.
@@
Up to a factor $(1+o(1)$ we have $\Delta N$ possibilities for $U^{*},\tau^{*},B$ and $N$ possibilities for $U,\tau,B$ and the uniform distribution on both spaces. If follows from Lemma 4.9 that the probability of the produced four-tuple is at most $\Delta 2^{O(s)}(\log n/\Delta^{2})^{s}$ that of initial four-tuple.
@@
$2^{O(s)}(\Delta t)^{s}(\log n)^{2s}\Delta(\log n/\Delta^{2})^{s}\leq 2^{O(s)}((\log n)^{2}t/\Delta)^{s}.$
@@
If we allow $r$ holes then the number of possibilities of $U^{*}$ increases by a factor at most $n^{2r}$. The number of $\tau^{*}$ is different as some sums are no longer 0, but the sum at each mini-square is uniquely determined by $U^{*}$. Once this is fixed the number of $\tau^{*}$ is at most the number of almost complete matchings by Lemma 3.8.
For there to be a hole, by the balance condition, all the at least $C\log n/8$ edges between two super-squares must be present in $J^{*}$. This implies that $r=O(s/\log n)$ and the factor $n^{O(r)}$ can be absorbed in the factor $2^{O(s)}$. This completes the proof also in this case.
@@ 5 The lower bound for proof size
Theorem 5.1
Assume that $d\leq O(\frac{\log n}{\log{\log n}})$ and let $n$ be an odd integer. Then any depth-$d$ Frege proof of the functional onto PHP on the $n\times n$ grid requires total size
$\mbox{exp}(\Omega(n^{1/(2d-1)}(\log n)^{O(1)})).$
Proof:  The proof follows the standard path. We use a sequence of restrictions and after the $i$th restriction any sub-formula of the proof of original depth at most $i$ is in the range of the $t$-evaluation.
Assuming this, after $d$ restrictions, any sub-formula of the proof is in the range of the $t$-evaluation. By Lemma 3.3 what remains is a smaller functional onto PHP instance and by Lemma 2.14, provided the size of the remaining grid is significantly larger than $t$, the proof cannot derive contradiction. Let us give some details.
In the base case of depth 0, each literal naturally gives a 1-evaluation. Now suppose that size (and hence the number of sub-formulas) of the proof is $2^{S}$.
@@
Indeed, consider Lemma 4.1 and take constants large enough so that the failure probability for an individual formula is bounded by $2^{-2S}$. By the union bound, except with probability $2^{-S}$, at step $i$ all formulas originally of depth $i$ are in the range of of a $2S$ evaluation.
@@
such a proof cannot refute the PHP. This concludes the proof of the theorem.
@@ 6 Multi-switching
The extension to multi-switching only requires minor modifications and we start by stating the lemma.
Lemma 6.1
There is a constants $A$ such that the following holds. Consider formulas $F_{i}^{m}$, for $m\in[M]$ and $i\in[n_{m}]$, each associated with a decision tree of depth at most $t$ and let $F^{m}=\lor_{i=1}^{n_{m}}F_{i}^{m}$. Let $\sigma$ be a random full restriction from the space of restrictions defined in Section 3. Then the probability that $(F^{m})_{m=1}^{M}$ cannot be represented by an $\ell$ common partial decision tree of depth at most $4s$ is at most
$\Delta M^{4s/\ell}\big{(}A(\log n)^{7}t\Delta^{-1}\big{)}^{s}.$
Proof:  As in the proof of the single switching lemma we outline a process to construct an $\ell$ common decision tree. We prove that this process is unlikely to ask more than $4s$ variables in a similar way to the single switching lemma.
Let us define the process of constructing a $\ell$ common decision tree. We treat the formulas for increasing values of $m$. We have counter $j$, starting at 0, indicating the number of times we have found a long branch in a decision tree. We have an information set $I^{+}$ which initially is empty.
@@
After this branch has been discovered we ask, now in the $\ell$ common decision tree, for the partner of any chosen element in $I^{j*}=\cup_{i=1}^{g_{j}}I_{i}^{j}$. The answers jointly with the matching pairs from $\pi_{2}$ in $I^{j*}$ form a new information set $I^{j*+}$. This information set is added to $I^{+}$ and we look for the next formula that needs a decision tree of depth at least $\ell$. It might be the same formula $F^{m_{j}}$, but the situation is new as we an updated set $I^{+}$ and we know that the previous forceable branch is now not followed.
@@
Proof:  The parts from $\pi_{2}$ are obviously disjoint and we need only consider the chosen mini-squares. For the same value of $j$, this lemma follows by the same argument as Lemma 4.5. As $I^{j*+}$ contains each $J^{j}_{i}$ and sets $J_{i}^{j^{\prime}}$ for $j^{\prime}>j$ are disjoint from $I^{j*+}$ the general case is also true.
@@
A difference compared to the single switching lemma is the sets $I^{*j+}$. Again the information pieces from $\pi_{2}$ do not give a problem as they never change. For each chosen mini-square in $I_{i}^{j}$ we need to specify a partner inside $I^{*j+}$. This may or may not belong to some $I_{i^{\prime}}^{j}$. If it does we can specify this at cost $O(1)$ but if not we can specify it at cost $O(\log n)$ as it is alive in $\rho^{*}$.
For each information piece in $J_{i}^{j}$ we thus have, in the worst case, to discover one mini-square at cost $t$, one at cost $\Delta$ and then six mini-squares at cost $O(\log n)$. This are two partners in $I_{i}^{j}$ and then up to four partners in $I^{*j+}$. This implies that the total external cost at stage $j$ is at most $2^{O(s_{j})}(\log n)^{6s_{j}}(t\Delta)^{s_{j}}$.
As any edge in $J_{i}^{j}$ might result in four queries in the common decision tree, if we have $4s$ questions in the this tree, then $\sum s_{j}\geq s$ and let us for convenience assume we have equality. The total cost of external information is in such a case $2^{O(s)}(\log n)^{6s}(t\Delta)^{s}$.
Lemma 4.9 applies also in this case from a similar calculation to that of the standard switching lemma we conclude that the total probability of quadruples $\tau$, $U$ $\pi_{2}$, $B$ that can reconstructed for a single choice of which formulas to process is bounded by
$M^{4s/\ell}\Delta 2^{O(s)}((\log n)^{7}t\Delta^{-1})^{s}.$
@@ 7 The lower bound for number of lines
Theorem 7.1
Assume that $d\leq O(\frac{\log n}{\log{\log n}})$ and let $n$ be an odd integer and $M$ a parameter. Then any depth-$d$ Frege proof of the functional and onto PHP on the $n\times n$ grid where each line is of size $M$ requires at least
@@
Proof:  Given that we established the multi-switching lemma the situation is now essentially the same as in [11].
@@
In our first application we set $4s_{1}=\log N$ and we conclude, by the union bound, that each line with high probability admits its own own $s_{1}$-common $\ell$-evaluation.
@@
Continuing similarly, in the $i$th iteration we use $s_{i}=2^{i-1}\log N$. The final common decision tree is of of depth at most $\sum_{j=0}^{i-1}s_{j}=O(2^{d}\log N)$.
After $d$ rounds of restrictions we have reduced the size of the PHP from $n$ to $n/(\log n)^{O(d)}(\log M)^{2d-2}$. If this is larger than $150(\sum_{j=0}^{d}s_{j}+\log M)$ we can conclude that there is no proof and this gives the bound of the theorem.
@@ 8 Final words
It seems that to be able to prove a switching lemma for a space of restrictions one essential property is that for any given variable, the probability of setting it to either constant should be significantly higher than keeping it undetermined. In the unrestricted PHP the probability that any variable is true is about $\frac{1}{n}$ and to make the probability of the variables remaining undetermined smaller we must go from size $n$ to a size smaller than $\sqrt{n}$. With such a quick decrease in size we can only apply the switching lemma $O(\log\log n)$ times. In the graph PHP on the grid the probability of a variable being true to about $\frac{1}{4}$ and at the same time the probability of keeping a variable undetermined is about $\log n/\Delta$. It is harder to preserve a graph PHP, but this is made possible by using augmenting paths as the new variables. Thus it seems that both decision were forced upon us but certainly there might be other possibilities and it would be interesting to see alternative proofs for a switching lemma for a space of restrictions that preserve the PHP.

### v3

@@ 4 The switching lemma
Both $\tau$ and $\tau^{*}$ are essentially almost complete matchings. We do not allow $\tau$ to be lopsided but this only reduces the number of possibilities, by Lemma 3.9, by a factor $1+o(1)$. It is the case that $\tau^{*}$ is only slightly more general but we only need an upper bound on the number of possibilities. On the other hand, if $U^{*}$ does not contain the designated survivor the sum of all adjacent $y_{P}$ (in $\pm 1$-notation) around this mini-square is not zero. The number of $\tau^{*}$ with these fixed sums is, however, smaller by Lemma 3.11. We conclude that the number of $\tau^{*}$ is bounded by $(1+o(1))$ times the number of different $\tau$.
There is no difference between $B$ and $B^{*}$ and thus the number of alternatives is the same. We conclude that if we denote the number of different $(\tau,U,B)$ by $N$, then the number of triples $(\tau^{*},U^{*},B^{*})$ is bounded by $\Delta N(1+o(1))$.
The big difference is in the probability of picking the restriction $\pi_{2}$ and $\pi_{2}^{*}$. The reason is that $\pi_{2}$ contains $k$ pairs and $\pi_{2}^{*}$ only ksk-s pairs. It is true that $\pi_{2}$ is also not unbalanced, but as this, by Lemma 3.13, only changes the number by a factor $(1+o(1))$ we ignore this here and absorb this factor in the error term. We do have that both $\pi_{2}$ and $\pi_{2}^{*}$ are matchings and thus we have to be slightly careful as the number of ways to extend a matching by one more pair depends on the current matching. In order to do this we put a probability distribution also on $\pi_{2}^{*}$. The random process to pick it is as the one picking $\pi_{2}$ (described on page 3.4) but only with fewer pairs picked and we ignore the balance conditions.
@@
Let $D$ denote the total number of possible pairs of mini-squares. We have that $D=\Delta^{2}(n/T)^{2}(2+o(1))$. This follow as there are $\Delta(n/T)^{2}$ mini-squares. Most of them (except at the perimeter) have $4\Delta$ possible partners and hence the number of pairs is $(1+o(1))\Delta(n/T)^{2}4\Delta/2$. We have the following lemma.
Lemma 4.18.
@@
Suppose $\pi_{2}$ is of size $k$ and $\pi_{2}^{*}\subset\pi_{2}$ is of size ksk-s and assume that $\Delta=\omega(\log n)$. Then $Pr[\pi_{2}]\leq 2^{o(s)}(k/D)^{s}Pr[\pi_{2}^{*}]$ which in its turn is bounded by $2^{O(s)}(\log n/\Delta^{2})^{s}Pr[\pi_{2}^{*}]$.
Proof 4.19.
@@
The way we define the probability space of partial matchings is slightly unusual and we need to be careful. There are $k!$ different ways to pick $\pi_{2}$. The probability of picking $\pi_{2}$ in a particular order $\gamma$ is $\prod_{i=1}^{k}t_{i,\gamma}^{-1}$ where $t_{i,\gamma}$ is the number of pairs of mini-squares available at stage $i$ assuming the order $\gamma$. As $\Delta=\omega(\log n)$ we have that $k$ is $o(\Delta(n/T)^{2})$. In other words very few of the mini-squares are picked at any stage of the process and hence $t_{i,\gamma}$ is on the form (1o(1))D(1-o(1))D but we mostly use much more detailed bounds below.
Similarly, the probability of $\pi_{2}^{*}$ can be written as
@@
where $\gamma^{*}$ determines the order in which its ksk-s pairs are added. Each picked pair of mini-squares eliminates at most $8\Delta$ possible other pairs and hence each $t_{i,\gamma}$ and $t_{i^{*},\gamma^{*}}$ is at least $D-8k\Delta$. The product giving the probabilities do have many factors so we need more precise information.
Say that $\gamma$ and $\gamma^{*}$ are associated (written as $\gamma\sim\gamma^{*}$) if the ksk-s common elements are in the same order. In such a situation for each $i^{*}$ there is an $i$ corresponding to the same element such that $t_{i^{*},\gamma^{*}}\geq t_{i,\gamma}\geq t_{i^{*},\gamma^{*}}-8s\Delta$. Restricting the product to these paired indices and using $t_{i,\gamma}\geq D/2$ we have
@@
$(1-16s\Delta/D)^{k}=2^{-\Theta((s\log n)/\Delta)}=2^{-o(s)},$
@@
giving the first bound of the lemma. The second bound is obtained using the definitions of $k$ and $D$.
@@
Up to a factor $(1+o(1))$ we have $\Delta N$ possibilities for $(U^{*},\tau^{*},B^{*})$ and $N$ possibilities for $(U,\tau,B)$ where we have the uniform distribution on both spaces. It follows from Lemma 4.18 that the probability of the produced four-tuple is at most $\Delta 2^{O(s)}(\log n/\Delta^{2})^{s}$ that of initial stared four-tuple.
@@
$2^{O(s)}(\Delta t)^{s}(\log n)^{2s}\Delta(\log n/\Delta^{2})^{s}\leq\Delta 2^{O(s)}((\log n)^{3}t/\Delta)^{s}.$
@@
If we allow $r$ holes then the number of possibilities of $U^{*}$ increases by a factor at most $n^{2r}$. The number of $\tau^{*}$ is different as some sums are no longer 0, but the sum at each mini-square is uniquely determined by $U^{*}$. Once this is fixed, the number of $\tau^{*}$ is at most the number of almost complete matchings by Lemma 3.11.
For there to be a hole, by the balance condition, all the at least $C\log n/8$ edges between two super-squares must be present in $J^{*}$. This implies that $r=O(s/\log n)$ and the factor $n^{O(r)}$ can be absorbed in the factor $2^{O(s)}$.
Another difference is that if we have $r$ holes then $\pi_{2}^{*}$ contains k+r/2sk+r/2-s pairs. Also the factor resulting from this change can be absorbed in the $2^{O(s)}$ factor. This completes the proof also in case when there are holes.
@@ 5 The lower bound for proof size
Theorem 5.1.
Assume that $d\leq O(\frac{\log n}{\log{\log n}})$ and let $n$ be an odd integer. Then any depth-$d$ Frege proof of the functional onto $\operatorname{PHP}$ on the $n\times n$ grid requires total size
$\exp(\Omega(n^{1/(2d-1)}(\log n)^{O(1)})).$
Proof 5.2.
The proof follows the standard path. We use a sequence of restrictions and after the $i$th restriction any sub-formula of the proof of original depth at most $i$ is in the range of the $t$-evaluation.
Assuming this, after $d$ restrictions, any sub-formula of the proof is in the range of the $t$-evaluation. By Lemma 3.4 what remains is a smaller functional onto $\operatorname{PHP}$ instance and by Lemma 2.9, provided the size of the remaining grid is significantly larger than $t$, the proof cannot derive contradiction. Let us give some details.
In the base case of depth 0, each literal naturally gives a $1$-evaluation. Now suppose that the size (and hence the number of sub-formulas) of the proof is $2^{S}$.
@@
Indeed, consider Lemma 4.1 and take constants large enough so that the failure probability for an individual formula is bounded by $2^{-2S}$. By the union bound, except with probability $2^{-S}$, at step $i$ all formulas originally of depth $i$ are in the range of a $2S$ evaluation.
@@
such a proof cannot refute the $\operatorname{PHP}$ and we get a contradiction. Rearranging we get the bound for $S$ claimed in the theorem. This concludes the proof.
@@ 6 Multi-switching
The extension to multi-switching only requires minor modifications and we start by stating the important lemma.
Lemma 6.1.
There is a constant, $A$, such that the following holds. Consider formulas $F_{i}^{m}$, for $m\in[M]$ and $i\in[n_{m}]$, each associated with a decision tree of depth at most $t$ and let $F^{m}=\lor_{i=1}^{n_{m}}F_{i}^{m}$. Let $\sigma$ be a random full restriction from the space of restrictions defined in Section 3. Then the probability that $(F^{m})_{m=1}^{M}$ cannot be represented by an $\ell$ common partial decision tree of depth at most $4s$ is at most
$\Delta M^{4s/\ell}\big(A(\log n)^{6}t\Delta^{-1}\big)^{s}.$
Proof 6.2.
As in the proof of the single switching lemma we outline a process to construct an $\ell$ common decision tree. We prove that this process is unlikely to ask more than $4s$ variables in a similar way to the single switching lemma.
Let us define the process of constructing an $\ell$ common decision tree. We treat the formulas for increasing values of $m$. We have counter $j$, starting at 0, indicating the number of times we have found a long branch in a decision tree. We have an information set $I^{+}$ which initially is empty.
@@
After this branch has been discovered we ask, now in the $\ell$ common decision tree, for the partner of any chosen element in $I^{j*}=\cup_{i=1}^{g_{j}}I_{i}^{j}$. The answers jointly with the matching pairs from $\pi_{2}$ in $I^{j*}$ form a new information set $I^{j*+}$. This information set is added to $I^{+}$ and we look for the next formula that needs a decision tree of depth at least $\ell$. It might be the same formula $F^{m_{j}}$, but the situation is new as we have an updated set $I^{+}$.
@@
Proof 6.4.
The parts from $\pi_{2}$ are obviously disjoint and we need only consider the chosen mini-squares. For the same value of $j$, this lemma follows by the same argument as Lemma 4.11. As $I^{j*+}$ contains each $J^{j}_{i}$ and sets $J_{i}^{j^{\prime}}$ for $j^{\prime}>j$ are disjoint from $I^{j*+}$ the general case is also true.
@@
A difference compared to the single switching lemma is the sets $I^{*j+}$. Again the information pieces from $\pi_{2}$ do not give a problem as they never change. For each chosen mini-square in $I_{i}^{j}$ we need to specify a partner inside $I^{*j+}$. This may or may not belong to some $I_{i^{\prime}}^{j}$. If it does, we can specify this at cost $O(1)$ but if not we can specify it at cost $O(1)+\log\log n$ bits as it is alive in $\rho^{*}$.
For each information piece in $J_{i}^{j}$ we thus have, in the worst case, to discover one mini-square at cost $t$, one at cost $\Delta$ and then six mini-squares at cost $O(1)+\log\log n$ bits. These are two partners in $I_{i}^{j}$ and up to four partners in $I^{*j+}$. This implies that the total external cost at stage $j$ is at most $s_{j}(O(1)+6\log\log n+\log t+\log{\Delta})$ bits.
As any edge in $J_{i}^{j}$ might result in four queries in the common decision tree, if we have $4s$ questions in this tree, then $\sum s_{j}\geq s$ and let us for convenience assume we have equality. The total cost of external information is in such a case $2^{O(s)}(\log n)^{6s}(t\Delta)^{s}$.
Lemma 4.18 applies also in the current case and from a similar calculation to that of the standard switching lemma, we conclude that the total probability of quadruples $\tau$, $U$ $\pi_{2}$, $B$ that can be reconstructed is bounded by
$M^{4s/\ell}\Delta 2^{O(s)}((\log n)^{6}t\Delta^{-1})^{s}.$
@@ 7 The lower bound for number of lines
Theorem 7.1.
Assume that $d\leq O(\frac{\log n}{\log{\log n}})$ and let $n$ be an odd integer and $M$ a parameter. Then any depth-$d$ Frege proof of the functional and onto $\operatorname{PHP}$ on the $n\times n$ grid where each line is of size $M$ requires at least
@@
Proof 7.2.
Given that we established the multi-switching lemma the situation is now essentially the same as in [12].
@@
In our first application we set $4s_{1}=\log N$ and we conclude, by the union bound, that each line with high probability admits its own $s_{1}$-common $\ell$-evaluation.
@@
Continuing similarly, in the $i$th iteration we use $s_{i}=2^{i-1}\log N$. The final common decision tree is of depth at most $\sum_{j=0}^{i-1}s_{j}=O(2^{d}\log N)$.
After $d$ rounds of restrictions we have reduced the side length of the grid on which we define the $\operatorname{PHP}$ from $n$ to $n/(\log n)^{O(d)}(\log M)^{2d-2}$. If this is larger than $300(\sum_{j=0}^{d}s_{j}+\log M)$ we can conclude that there is no proof and this gives the bound of the theorem.
@@ 8 Final words
It seems that to be able to prove a switching lemma for a space of restrictions one essential property is that for any given variable, the probability of setting it to either constant should be significantly higher than keeping it undetermined. In the unrestricted $\operatorname{PHP}$ the probability that any variable is true is about $\frac{1}{n}$ and to make the probability of the variables remaining undetermined smaller we must go from size $n$ to a size smaller than $\sqrt{n}$. With such a quick decrease in size we can only apply the switching lemma $O(\log\log n)$ times. In the graph $\operatorname{PHP}$ on the grid the probability of a variable being true to about $\frac{1}{4}$ and at the same time the probability of keeping a variable undetermined is about $\log n/\Delta$. It is harder to preserve a graph $\operatorname{PHP}$, but this is made possible by using augmenting paths as the new variables. Thus it seems that both decisions were forced upon us but certainly there might be other possibilities and it would be interesting to see alternative proofs for a switching lemma for a space of restrictions that preserve the $\operatorname{PHP}$.